SDC Travel Budget

Travel budget was based on a “bottoms-up” estimate from each institution

• scrutinised for “reality”

• required to be consistent with mission

Total travel budget requested for 1993-94 is $327.4K

• Based on 8.5 FTE eligible NSERC grant holders

• Includes travel for 8.5 FTE collaborators

Estimated travel budget for 1992-93 is ∼ $130K

• Based on 6.5 FTE eligible NSERC grant holders

• Includes travel for 0.5 FTE collaborators